# Spin-glass states generated in van der Waals magnet CrGeTe by alkali-ion intercalation

### v1

@@ Abstract
Tuning magnetic properties in layered van der Waals (vdW) materials has captured a significant attention due to the efficient control of ground-states by heterostructuring and external stimuli. Electron doping by electrostatic gating, interfacial charge transfer and intercalation is particularly effective in manipulating the exchange and spin-orbit properties, resulting in a control of Curie temperature () and magnetic anisotropy. Here, we discover an uncharted role of intercalation to generate magnetic frustration. As a model study, we intercalate Na atoms into the vdW gaps of pristine CrGeTe (CGT) where generated magnetic frustration leads to emerging spin-glass states coexisting with a ferromagnetic order. A series of dynamic magnetic susceptibility measurements/analysis confirms the formation of magnetic clusters representing slow dynamics with a distribution of relaxation times. The intercalation also modifies other macroscopic physical parameters including the significant enhancement of from 66K to 240 K and the switching of magnetic easy-hard axis direction. Our study identifies intercalation as a unique route to generate emerging frustrated spin states in simple vdW crystals.
@@ 1 Introduction
Among the vdW magnetic materials, the families of chromium halides (i.e. CrI, CrBr and CrCl), lamellar ternary chalcogenides (i.e. CrGeTe and CrSiTe), metal thiophosphates (MPS, M = Fe, Mn and Ni) and iron-based tellurides (i.e. FeGeTe and FeGeTe), have been intensively and extensively studied [18, 19, 20, 21, 22]. One common aspect shared by most of these materials is the magnetic ordering transition temperature, Curie temperature () for ferromagnets, lying below room temperature [14], with a few exceptions including FeGeTe [14] and FeGaTe [23]. In order to fully realise the potential of 2D magnetic materials for practical applications, a consequential drive is taking place in either discovering new materials with above-room-temperature , or to elevate of existing systems by external means. For the latter, taking CrGeTe (CGT) as an example, its of 66 K [24, 25] is reported to be significantly enhanced by electrostatic gating [26, 27], strain [28, 29] and intercalation [7].
@@
The modification of magnetic properties in vdW systems has been primarily focused on parameters such as and magnetic anisotropy constants that arise from the exchange and spin-orbit interaction respectively. Electron doping alters the type and strength of the exchange interaction [31, 32], e.g. via the change of the Fermi level position for itinerant ferromagnets [33], and the number of carriers that play a leading role in carrier-mediated indirect Ruderman–Kittel–Kasuya–Yosida (RKKY) and double-exchange mechanisms for less-conductive magnets [26]. Doped electrons populate in electronic bands with a specific orbital character, modifying the spin-orbit property impacting on magnetic anisotropies [34]. These material properties represent the homogeneous nature of systems. By contrast, some important yet unanswered questions in researching vdW magnetic systems are: what arise when a vdW magnetic system is introduced with inhomogeneous carrier-doping, and what are resultant parameter changes that we can experimentally access to? To address this with a model case, we employ Na-ion intercalation into the vdW gaps of pristine CrGeTe where Na ions are inhomogeneously intercalated within the gaps. We actively use this nature to intentionally create disorder and study its consequence to magnetism.
@@
While inhomogeneity might be considered negative since it breaks the periodicity/coherency of crystallinity in real space, it can however generate new physics that cannot be achieved otherwise. In magnetism, spin-glass states [35] emerge because of the presence of multiple local energy minima that the system freezes into, generated by inhomogeneity and frustration/competition between the ferromagnetic and anti-ferromagnetic exchange interactions. Hopping between the energy minima originates experimentally measurable magnetic responses that are time-, temperature- and magnetic-field-dependent [10]. Characteristic of the phase transition, the spin-glass has an order parameter defined by Edwards and Anderson [37], together with the critical temperature of its freezing, . The research field of spin-glasses is vast and diverse, ranging from canonical spin-glass states often observed in non-magnetic materials with randomly distributed magnetic impurities [38] to more complex spin-glassy behaviours in frustrated ferromagnets, called reentrant spin-glasses [39]. Spin glass formation in highly correlated systems is also known to herald yet more exotic behaviour such as high-temperature superconductivity in the cuprates [40]. The realisation of glassy magnetic states in vdW materials therefore suggests their significant potential for supporting new magnetic and electronic states, formed by tuning material properties via substitution, intercalation and exfoliation or heterostructuring.
@@ 2 Results Discussion
2.1 Characterisation of Na-CrGeTe
Figure 1 (a b) depicts the CGT and Na-CGT systems, respectively, and illustrates the presence of sodium-ions in the vdW gaps. Single CGT crystals were intercalated with sodium ions via the ammonia-based method [41] (details in Methods section). Structural changes in the intercalated CGT are characterised using X-ray diffraction (XRD) and Raman spectroscopy. XRD results in Fig. 1 (c d) show that the intercalated sample preserves the original crystalline structure of CGT. The c-axis parameter of pristine CGT is 6.8 Å, which corresponds to the distance between the centre of adjacent layers. A downshift is observed for all 00l-peaks after intercalation, indicating an increase in the interlayer spacing by 0.04 Å. This small increase in interlayer spacing upon intercalation is also observed in other intercalated compounds [42]. Raman spectroscopy is also performed for a large number of spatial positions on the sample surface, and the peak structure is maintained, consistent with the indication from XRD. However, quantitative analysis of the three most significant Raman modes (, , and ) shows that there are statistically significant changes to phonon energy and lifetime with intercalation. There are no major changes to the three main Raman modes that arise from intralayer vibrations of CGT, as expected, as intercalation should not change the bonding within the layers themselves. The small changes in peak energy and width (which is determined by lifetime) are consistent with carrier doping [9, 8] (see Supplementary Information Fig.S2-S3).
@@ 2.1 Characterisation of Na-CrGeTe
Figure 2 (a) shows temperature-dependent magnetisation along the ab-plane orientation in both CGT and Na-CGT, and demonstrates the impact of Na intercalation on magnetisation in the static regime. In contrast to of CGT 66 K [24], Na-CGT shows its magnetisation values maintained to a higher temperature up to , implying this as its . Similar enhancement in CGT has been reported by organic molecule intercalation (to 208 K) [7] and by pressure (to 250 K) [28] where in all cases metallic transport accompanies the enhancement. Since scales with the exchange coupling strength, it is possible to interpret in our sample that doped electrons enhance the exchange coupling strength and at the same time these less-localised electrons can contribute to transport properties. In pristine CGT, electrons are fully localised and the super-exchange mechanism that arises from virtual electron excitations between two neighbouring Cr ions via a Te ion with the 90 Kanamori-Goodenough mechanism [46]. Local Hund’s rules within both Te and Cr ions stabilise a ferromagnetic order. Through this process, electrons have to be virtually excited into the conduction band via the bandgap, the exchange energy gain by this process is marginal, hence low in CGT. Doping real electrons to Cr ions changes their ionic states from Cr into Cr, activating the double-exchange mechanism [47] that bolsters the exchange interaction. It may also be relevant that since Cr is a Jahn-Teller ion, its substitution for Cr will lead to local distortions and increased local anisotropy that potentially enhances any spin glass behaviour that we discuss below.
CGT is known to have uniaxial magnetic anisotropy with the out-of-plane easy axis [24, 25]. Magnetisation measurements as a function of magnetic field (- curves) are performed for Na-CGT and the results at 10 K for the ab-plane and c-axis orientations are shown in Fig. 2 (b). It is evident that after intercalation, spins prefer to align within the ab-plane compared to the c-axis in CGT, implying the sign change of the uniaxial anisotropy term upon intercalation. Figure 2 (c) shows the - curves of Na-CGT at 200 K, showing a clear magnetic response for both orientations with the sustained uniaxial anisotropy. As a proof of the long-range magnetic order, such a response disappears at 295 K where only a linear - trend is observed and shown in Fig.S5(d) Supplementary Information.
@@ 2.2 Emergent spin-glass nature
Another striking property which emerges via the intercalation process is spin-glass like behaviour in Na-CGT, a feature absent in pristine CGT. It is important to note that spin-glass states have not been experimentally reported in other heavily studied 2D magnetic systems [19, 20], such as CrI, CrBr, FeGeTe, as well as CGT with carrier doping by organic molecules [7] and electrostatic gating [26, 27], and pressure application [28, 29].
Spin-glass behaviour in Na-CGT is identified as the divergence of magnetisation in temperature between the zero-field cooled (ZFC) and field-cooled (FC) protocols, at small magnetic fields. As shown in Fig. 3 (a), a bifurcation between the two cases is observed below a temperature at which a steep increase in magnetisation takes places. The magnetisation of the ZFC curve shows a defined broad peak, whereas in the FC protocol, the magnetisation keeps on increasing as the sample is further cooled. The local maximum is observed and we define this as [48, 49]. This bifurcation in the magnetisation is minimised as the size of magnetic field in increased (see Supplementary Information). Such splitting and irreversibility in the magnetic memory of the system are recognised as characteristic features which can arise from various phenomena, including spin-liquid, super-paramagnetic, disordered anti-ferromagnetic and spin or cluster glass states [10, 38].
@@
The observed spin-glass property has formed within the ferromagnetically ordered matrix background. This particular type of spin freezing below a ferromagnetic transition is referred to as the reentrant spin glass state [39]. Here, the fluctuation of local exchange interaction is sufficient enough to produce the local energy minima that stabilise frozen spin-glass states against the ferromagnetic order [52, 53]. Phenomenologically, this condition is met when the averaged exchange coupling strength for the th and th site spins <> and its local fluctuation average <> are comparable to each other [39], where < · · · > denotes an average over the entire spin system. In Na-CGT, this situation is likely established due to inhomogeneous Na intercalation that induces an electron-doping distribution. Since the exchange coupling strength is a function of carrier density between magnetic ions, such doping inhomogeneity brings about the fluctuation and hence local competition of the exchange interaction, a key ingredient for the spin-glass states. While such atomic scale fluctuation is not directly measured in our study, the clear spin-glass signature observed in our study strongly indicates its presence.
@@
For more quantitative analysis, a series of analytical procedures is carried out with phenomenological scaling and power-law models [10, 54] to further inquest and interpret the data with in as a function of frequency. The Mydosh parameter [10], , is measured as the respective variation of peak-maximum position per frequency, , defined as, . Here, corresponds to the difference in experimental temperature values of peak-maximum at high and low frequencies; is the difference in excitation frequencies for the given value. K sets a criterion for differentiating between superparamagetic transition () and spin-glass behaviour () [10, 55, 56, 57], and a similar frequency-dependent shift in has not been reported for anti-ferromagnetic or ferromagnetic phases at low -values [10]. For Na-CGT, is experimentally determined to be 0.01 which clearly suggests its spin-glass nature, sitting between the boundary of canonical and cluster spin glasses[58, 59].
@@
In the next step, the standard theory of dynamic scaling near the phase transition temperature for given excitation frequencies is employed [10, 54]. This theory relies on a power law dependence of relaxation time (experimental attempt time), , as , where is the characteristic relaxation time related to the microscopic flipping time for dynamical fluctuating entities, such as a single moment or a cluster. corresponds to the freezing temperature in the limit of , and is the critical exponent. The power law equation is simplified into a linear form using a logarithmic base, and is adjusted in order to get the best linearity [54], where the goodness of fit is determined from the highest -coefficient value. Figure 3 (c) shows the result of this fitting process of the power law. is found to be 33.7 K, and then using this value, a linear least-squares fitting method is utilised to extrapolate s and . The obtained value of for Na-CGT lies in the range ( to s) typical for cluster spin-glass systems [59, 60] compared to canonical spin glasses [61, 62] ( to s) where the fluctuation of individuals moments happens at a faster rate. This indicates slower spin dynamics in Na-CGT at the transition temperature and it is most likely due to the presence of interacting spin-clusters in the system instead of individual spins. The value of zv for typical systems with spin-glass behaviour is reported between 4 and 12 [59, 62], and indeed this is also the case for Na-CGT, providing another experimental evidence for the cluster spin-glass nature in Na-CGT.
@@
Finally, the Casimir-du Pré equations are utilised [10, 12], in order to describe the dependence of both real and imaginary parts of the ac susceptibility on the excitation frequency at a given temperature. Figures 4 (a) and (b) show good agreement between the experimental data for both real/imaginary parts and the modified equations (Equations S3 S4) for different temperatures. The slow dynamics, arising in the vicinity of 35 K, are described by the generalised Debye model that takes into account the entities (i.e. spin-clusters for Na-CGT) having finite interactions leading to a system having a distribution of relaxation times due to the presence of cooperative effects [12]. is plotted against at a single temperature value and for different frequencies as shown in Fig. 4 (c), and we fit them by the following Cole-Cole equation:
where represents the susceptibility in the limit of zero frequency, known as isothermal susceptibility; is the adiabatic susceptibility for the high frequency limit; is a dimensionless parameter to represent its relaxation nature with 0 being the case for single relaxation time and 1 being for a broad distribution of relaxation times. The fact that the fit curves show an arc centered away from the origin indicates the presence of a distribution of relaxation times, compared to a semi-circle appearing for the system with a single relaxation time [12]. The presence of such distribution can be attributed to magnetic clusters with different sizes and anisotropies, each experiencing different relaxation times in the whole system. Further quantitative analysis can be available in Supplementary Information. This series of analysis confirms that at low temperatures (in the vicinity of 35 K), intercalation of CGT by Na-atoms leads to the co-existence of a cluster spin-glass transition with a background ferromagnetic state. The analysis point towards a progressive freezing of spin clusters and slower dynamics.

### v3

# Spin-glass states generated in a van der Waals magnet by alkali-ion intercalation

@@ Abstract
Tuning magnetic properties in layered van der Waals (vdW) materials has captured a significant attention due to the efficient control of ground-states by heterostructuring and external stimuli. Electron doping by electrostatic gating, interfacial charge transfer and intercalation is particularly effective in manipulating the exchange and spin-orbit properties, resulting in a control of Curie temperature () and magnetic anisotropy. Here, we discover an uncharted role of intercalation to generate magnetic frustration. As a model study, we intercalate Na atoms into the vdW gaps of pristine Cr2Ge2Te6 (CGT) where generated magnetic frustration leads to emerging spin-glass states coexisting with a ferromagnetic order. A series of dynamic magnetic susceptibility measurements/analysis confirms the formation of magnetic clusters representing slow dynamics with a distribution of relaxation times. The intercalation also modifies other macroscopic physical parameters including the significant enhancement of from 66 K to 240 K and the switching of magnetic easy-hard axis direction. Our study identifies intercalation as a unique route to generate emerging frustrated spin states in simple vdW crystals.
@@ 1 Introduction
Among the vdW magnetic materials, the families of chromium halides (i.e. CrI3, CrBr3 and CrCl3), lamellar ternary chalcogenides (i.e. Cr2Ge2Te6 and Cr2Si2Te6), metal thiophosphates (MPS3, M = Fe, Mn and Ni) and iron-based tellurides (i.e. Fe3GeTe2 and Fe5GeTe2), have been intensively and extensively studied [18, 19, 20, 21, 22]. One common aspect shared by most of these materials is the magnetic ordering transition temperature, Curie temperature () for ferromagnets, lying below room temperature [14], with a few exceptions including Fe5GeTe2 [14] and Fe3GaTe2 [23]. In order to fully realise the potential of 2D magnetic materials for practical applications, a consequential drive is taking place in either discovering new materials with above-room-temperature , or to elevate of existing systems by external means. For the latter, taking Cr2Ge2Te6 (CGT) as an example, its of 66 K [24, 25] is reported to be significantly enhanced by electrostatic gating [26, 27], strain [28, 29] and intercalation [30].
@@
The modification of magnetic properties in vdW systems has been primarily focused on parameters such as and magnetic anisotropy constants that arise from the exchange and spin-orbit interaction respectively. Electron doping alters the type and strength of the exchange interaction [31, 32], e.g. via the change of the Fermi level position for itinerant ferromagnets [33], and the number of carriers that play a leading role in carrier-mediated indirect Ruderman–Kittel–Kasuya–Yosida (RKKY) and double-exchange mechanisms for less-conductive magnets [26]. Doped electrons populate in electronic bands with a specific orbital character, modifying the spin-orbit property impacting on magnetic anisotropies [34]. These material properties represent the homogeneous nature of systems. By contrast, some important yet unanswered questions in researching vdW magnetic systems are: what arise when a vdW magnetic system is introduced with inhomogeneous carrier-doping, and what are resultant parameter changes that we can experimentally access to? To address this with a model case, we employ Na-ion intercalation into the vdW gaps of pristine Cr2Ge2Te6 where Na ions are inhomogeneously intercalated within the gaps. We actively use this nature to intentionally create disorder and study its consequence to magnetism.
@@
While inhomogeneity might be considered negative since it breaks the periodicity/coherency of crystallinity in real space, it can however generate new physics that cannot be achieved otherwise. In magnetism, spin-glass states [35] emerge because of the presence of multiple local energy minima that the system freezes into, generated by inhomogeneity and frustration/competition between the ferromagnetic and anti-ferromagnetic exchange interactions. Hopping between the energy minima originates experimentally measurable magnetic responses that are time-, temperature- and magnetic-field-dependent [36]. Characteristic of the phase transition, the spin-glass has an order parameter defined by Edwards and Anderson [37], together with the critical temperature of its freezing, . The research field of spin-glasses is vast and diverse, ranging from canonical spin-glass states often observed in non-magnetic materials with randomly distributed magnetic impurities [38] to more complex spin-glassy behaviours in frustrated ferromagnets, called reentrant spin-glasses [39]. Spin glass formation in highly correlated systems is also known to herald yet more exotic behaviour such as high-temperature superconductivity in the cuprates [40]. The realisation of glassy magnetic states in vdW materials therefore suggests their significant potential for supporting new magnetic and electronic states, formed by tuning material properties via substitution, intercalation and exfoliation or heterostructuring. However, there are a limited number of reported studies on spin-glass states in vdW systems and all of them use magnetic disorders and frustrations arising in their host crystals [41, 42, 43, 44].
@@ 2 Results Discussion
2.1 Characterisation of Na-Cr2Ge2Te6
Figure 1 (a b) depicts the CGT and Na-CGT systems, respectively, and illustrates the presence of sodium-ions in the vdW gaps. Single CGT crystals were intercalated with sodium ions via the ammonia-based method [45] (details in Methods section). Structural and chemical changes in the intercalated CGT are characterised using X-ray diffraction (XRD), Raman spectroscopy and energy-dispersive X-ray (EDX) spectroscopy (see Supplementary Information Section 3). XRD results in Fig. 1 (c d) show that the intercalated sample preserves the original crystalline structure of CGT. The c-axis parameter of pristine CGT is 6.8 Å, which corresponds to the distance between the centre of adjacent layers. A downshift is observed for all 00l-peaks after intercalation, indicating an increase in the interlayer spacing by 0.04 Å. This small increase in interlayer spacing upon intercalation is also observed in other intercalated compounds [46]. Raman spectroscopy is also performed for a large number of spatial positions on the sample surface, and the peak structure is maintained, consistent with the indication from XRD. However, quantitative analysis of the three most significant Raman modes (, , and ) shows that there are statistically significant changes to phonon energy and lifetime with intercalation. There are no major changes to the three main Raman modes that arise from intralayer vibrations of CGT, as expected, as intercalation should not change the bonding within the layers themselves. The small changes in peak energy and width (which is determined by lifetime) are consistent with carrier doping [47, 48] (see Supplementary Information Fig.S2-S3).
@@ 2.1 Characterisation of Na-Cr2Ge2Te6
Figure 2 (a) shows temperature-dependent magnetisation along the ab-plane orientation in both CGT and Na-CGT, and demonstrates the impact of Na intercalation on magnetisation in the static regime. In contrast to of CGT 66 K [24], Na-CGT shows its magnetisation values maintained to a higher temperature up to , implying this as its . Our theoretical modeling also reproduces the trend of enhancement with the introduction of Na intercalated atoms in CGT (see Supplementary Information Fig.S14). Details of the Na distribution between the layers, complexity of the spin interactions and potential electron localisation are factors for such simulations. Experimentally, enhancement of in CGT has been reported by organic molecule intercalation (to 208 K) [30] and by pressure (to 250 K) [28] where in all cases metallic transport accompanies the enhancement. Since scales with the exchange coupling strength, it is possible to interpret in our sample that doped electrons enhance the exchange coupling strength and at the same time these less-localised electrons can contribute to transport properties. In pristine CGT, electrons are fully localised and the super-exchange mechanism that arises from virtual electron excitations between two neighbouring Cr ions via a Te ion with the 90∘ Kanamori-Goodenough mechanism [50]. Local Hund’s rules within both Te and Cr ions stabilise a ferromagnetic order. Through this process, electrons have to be virtually excited into the conduction band via the bandgap, the exchange energy gain by this process is marginal, hence low in CGT. Doping real electrons to Cr ions changes their ionic states from Cr3+ into Cr2+, activating the double-exchange mechanism [51] that bolsters the exchange interaction. It may also be relevant that since Cr2+ is a Jahn-Teller ion, its substitution for Cr3+ will lead to local distortions [52] and increased local anisotropy that potentially enhances any spin glass behaviour that we discuss below.
CGT is known to have uniaxial magnetic anisotropy with the out-of-plane easy axis [24, 25]. Magnetisation measurements as a function of magnetic field (- curves) are performed for Na-CGT and the results at 10 K for the ab-plane and c-axis orientations are shown in Fig. 2 (b). It is evident that after intercalation, spins prefer to align within the ab-plane compared to the c-axis in CGT, implying the sign change of the uniaxial anisotropy term upon intercalation, which is reproduced by our simulations with and without Na intercalants in CGT (see Supplementary Table S5). Figure 2 (c) shows the - curves of Na-CGT at 200 K, showing a clear magnetic response for both orientations with the sustained uniaxial anisotropy. As a proof of the long-range magnetic order, such a response disappears at 295 K where only a linear - trend is observed and shown in Fig.S6(d) Supplementary Information.
@@ 2.2 Emergent spin-glass nature
Another striking property which emerges via the intercalation process is spin-glass like behaviour in Na-CGT, a feature absent in pristine CGT. Spin-glass behaviour in Na-CGT is identified as the divergence of magnetisation in temperature between the zero-field cooled (ZFC) and field-cooled (FC) protocols, at small magnetic fields. As shown in Fig. 3 (a), a bifurcation between the two cases is observed below a temperature at which a steep increase in magnetisation takes places. The magnetisation of the ZFC curve shows a defined broad peak, whereas in the FC protocol, the magnetisation keeps on increasing as the sample is further cooled. The local maximum is observed and we define this as [53, 54]. This bifurcation in the magnetisation is minimised as the size of magnetic field in increased (see Supplementary Information). Such splitting and irreversibility in the magnetic memory of the system are recognised as characteristic features which can arise from various phenomena, including spin-liquid, super-paramagnetic, disordered anti-ferromagnetic and spin or cluster glass states [36, 38].
@@
The observed spin-glass property has formed within the ferromagnetically ordered matrix background. This particular type of spin freezing below a ferromagnetic transition is referred to as the reentrant spin glass state [39]. Here, the fluctuation of local exchange interaction is sufficient enough to produce the local energy minima that stabilise frozen spin-glass states against the ferromagnetic order [57, 58]. Phenomenologically, this condition is met when the averaged exchange coupling strength for the th and th site spins <> and its local fluctuation average <>2 are comparable to each other [39], where < · · · > denotes an average over the entire spin system. In Na-CGT, this situation is likely established due to inhomogeneous Na intercalation that induces an electron-doping distribution. Since the exchange coupling strength is a function of carrier density in CGT [26, 30], such doping inhomogeneity brings about the fluctuation and hence local competition of the exchange interaction, a key ingredient for the spin-glass states. While such atomic scale fluctuation is not directly measured in our experiments, we performed ab-initio simulations to calculate the exchange coupling strengths with Na intercalation, as shown in SI Sec. S7 where we clearly observe the change of the strengths by intercalation. However, as discussed later, the observed spin-glass states are a magnetic cluster type, which is rather difficult to reproduce by atomic-level disorders as we simulated. We speculate that the observed spin-glass states in Na-CGT originate from rich and complex mechanisms, possibly involving other inhomogeneity arising from inhomogeneous doping, such as the spatial distribution of magnetic anisotropy [26, 30].
@@
For more quantitative analysis of experimental results, a series of analytical procedures is carried out with phenomenological scaling and power-law models [36, 59] to further inquest and interpret the data with in as a function of frequency. The Mydosh parameter [36], , is measured as the respective variation of peak-maximum position per frequency, , defined as, . Here, corresponds to the difference in experimental temperature values of peak-maximum at high and low frequencies; is the difference in excitation frequencies for the given value. K sets a criterion for differentiating between superparamagetic transition () and spin-glass behaviour () [36, 60, 61, 62], and a similar frequency-dependent shift in has not been reported for anti-ferromagnetic or ferromagnetic phases at low -values [36]. For Na-CGT, is experimentally determined to be 0.01 which clearly suggests its spin-glass nature, sitting between the boundary of canonical and cluster spin glasses[63, 64].
In the next step, the standard theory of dynamic scaling near the phase transition temperature for given excitation frequencies is employed [36, 59]. This theory relies on a power law dependence of relaxation time (experimental attempt time), , as:
@@
where is the characteristic relaxation time related to the microscopic flipping time for dynamical fluctuating entities, such as a single moment or a cluster. corresponds to the freezing temperature in the limit of , and is the critical exponent. The power law equation is simplified into a linear form using a logarithmic base, and is adjusted in order to get the best linearity [59], where the goodness of fit is determined from the highest -coefficient value. Figure 3 (c) shows the result of this fitting process of the power law. is found to be 33.7 K, and then using this value, a linear least-squares fitting method is utilised to extrapolate s and . The obtained value of for Na-CGT lies in the range ( to s) typical for cluster spin-glass systems [64, 65] compared to canonical spin glasses [66, 67] ( to s) where the fluctuation of individuals moments happens at a faster rate. This indicates slower spin dynamics in Na-CGT at the transition temperature and it is most likely due to the presence of interacting spin-clusters in the system instead of individual spins. The value of zv for typical systems with spin-glass behaviour is reported between 4 and 12 [64, 67], and indeed this is also the case for Na-CGT, providing another experimental evidence for the cluster spin-glass nature in Na-CGT.
@@
Furthermore, the Casimir-du Pré equations are utilised [36, 55], in order to describe the dependence of both real and imaginary parts of the ac susceptibility on the excitation frequency at a given temperature. Figures 4 (a) and (b) show good agreement between the experimental data for both real/imaginary parts and the modified equations (Equations S3 S4) for different temperatures. The slow dynamics, arising in the vicinity of 35 K, are described by the generalised Debye model that takes into account the entities (i.e. spin-clusters for Na-CGT) having finite interactions leading to a system having a distribution of relaxation times due to the presence of cooperative effects [55]. is plotted against at a single temperature value and for different frequencies as shown in Fig. 4 (c), and we fit them by the following Cole-Cole equation:
where represents the susceptibility in the limit of zero frequency, known as isothermal susceptibility; is the adiabatic susceptibility for the high frequency limit; is a dimensionless parameter to represent its relaxation nature with 0 being the case for single relaxation time and 1 being for a broad distribution of relaxation times. The fact that the fit curves show an arc centered away from the origin indicates the presence of a distribution of relaxation times, compared to a semi-circle appearing for the system with a single relaxation time [55]. The presence of such distribution can be attributed to magnetic clusters with different sizes and anisotropies, each experiencing different relaxation times in the whole system. Further quantitative analysis are available in Supplementary Information. This series of analysis confirms that at low temperatures (in the vicinity of 35 K), intercalation of CGT by Na-atoms leads to the co-existence of a cluster spin-glass transition with a background ferromagnetic state. The analysis point towards a progressive freezing of spin clusters and slower dynamics.
Finally, we compare the spin-glass behaviour observed in Na-CGT and those in other vdW material systems reported so far [43, 41, 42, 69, 44, 70]. A critical difference between ours from the other vdW spin-glass systems is on the way to generate magnetic frustration. While our study demonstrates intercalation of non-magnetic atoms as a new route of generating magnetic frustration, other existing studies have used substitution of magnetic elements as a common method for creating spin-glass states. An example is 3d-transition metal phosphorus trisulfide XPS3 where X is substituted between two elements (such as Fe0.5Ni0.5 [43] and Mn0.5Fe0.5 [41, 42, 69]). There is a recent study by Wang et al. [70] where amorphous Cr-Ge-Te hosts spin glass behaviour and its driving mechanism is due to the presence of structural motifs corresponding to the known crystalline phases Cr-Te (n = 1, 2 and 3) with different Cr-Te-Cr bonding angles having a series of exchange interaction strengths. The importance of our study is that we maintain the original crystalline structure during the intercalation (as evident by XRD results) and generate spin-glass states without changing the Cr-Te-Cr bonding angles. It is worth highlighting that the relaxation rate extracted by the power law analysis (Eq.(1)) in the amorphous Cr-Ge-Te system is s which is one order faster than ours ( s) based on the same analysis. This difference in relaxation rates is consistent with our classification of the types of spin-glass state in amorphous Cr-Ge-Te (short-lived canonical spin-glass, with a small moment and rapid relaxation rate) and Na-CGT (cluster spin-glass with large moment and slower relaxation).